# Q-SLAM: Quadric Representations for Monocular SLAM

### v1

@@ 3.4 Joint Optimization of Rendering and Pose Estimation
During mapping, we map the 3D coordinates $\mathbf{x}$, and viewing directions $(\theta,\phi)$ to volume density, color and semantic logits $(\sigma,\textit{{c}},\textit{{s}})$. Subsequently, volume rendering is utilized to reconstruct color $\hat{\mathbf{C}}$, depth $\hat{\mathbf{D}}$, and semantic map $\hat{\mathbf{S}}$. For points $\{p_{i}|i=1,\cdots,M\}$ on a fitted quadric surface, the color and depth loss is calculated as follows:
@@
Here, the fitting error $\varepsilon$ functions as an uncertainty penalty. In cases where the fitting is not perfect, the quadric plane is less likely to exhibit similar texture and spatial features, resulting in reduced usefulness of the quadric transformer. Consequently, such patches are down-weighted by incorporating the fitting error $\varepsilon$. Following Zhi et. al. [50], we use a multi-class cross-entropy loss as the semantic loss $\mathcal{L}_{s}$. Hence the rendering loss is a weighted sum of $\mathcal{L}_{c}$, $\mathcal{L}_{d}$ and $\mathcal{L}_{s}$ with hyperparameters $\lambda_{1}$ and $\lambda_{2}$.
@@
For joint optimization, the camera pose $\mathbf{G}=\{g_{t}\}$, a $4\times 4$ transformation matrix, is converted to quaternion ($4\times 1$) and translation ($3\times 1$) vector, which are then taken as trainable parameters to the network, optimized together with the NeRF parameters while minimizing the rendering loss. In this way, the camera poses $\mathbf{G}$ and network parameters $\Theta$ can be jointly optimized, enhancing both the tracking and mapping accuracy.
@@ 4.1 Experimental Setup
Dataset. Q-SLAM is evaluated on a variety of datasets, including Replica [28], ScanNet [6], and TUM RGB-D [29] dataset. For evaluation of the reconstruction quality, we test our method on 8 synthetic scenes from Replica, which provides high-quality synthetic scenes, akin to the evaluation framework adopted by NeRF-SLAM [22]. Following GO-SLAM [49], we evaluate the tracking accuracy on ScanNet dataset which offers extensively annotated RGB-D scans of real-world scenarios, encompassing challenging short and long trajectories. Following Nice-SLAM [52], we also evaluate on various scenes on indoor TUM RGB-D dataset, with ground truth poses provided by a motion capture system. For camera tracking assessment, our approach is evaluated under two distinct modes: one utilizing ground truth and the other utilizing estimated depth from monocular images as inputs. The input images are resized to $640\times 320$ for Replica dataset, $320\times 240$ for ScanNet dataset and $512\times 384$ for TUM RGB-D dataset.
@@
Metrics.We evaluate tracking accuracy by aligning the estimated trajectory with the ground truth trajectory and computing the Root Mean Square Error (RMSE) of the Absolute Trajectory Error (ATE). In line with NeRF-SLAM [22], we employ the Depth L1 metric to evaluate depth prediction and utilize Peak Signal-to-Noise Ratio (PSNR), SSIM [39], and LPIPS [47] for image rendering evaluation.
Implementation Details. All experimental procedures are executed on an NVIDIA A6000 GPU using PyTorch 1.10.0. We use Adam as the optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The tracking module is adapted from Droid-SLAM [34], and pretrained weights are utilized to estimate depths and poses. Our mapping backbone is TensoRF [3], incorporating our proposed depth correction and quadric ray transformer. Diverging from methodologies where scene meshes are reconstructed using marching cubes on Signed Distance Function (SDF) values of queried points, Q-SLAM renders images and depths over the estimated camera trajectory and use TSDF-Fusion [46] for mesh construction with voxel size 1 cm. During the fitting process, only quadric surfaces with fitting error lower than the specified threshold undergo depth correction; otherwise, the uncorrected predicted depth is used for the supervision of NeRF training. For joint optimization, the scene representation parameters $\Theta$ are optimized for 5 steps, and the accumulated losses are then utilized to update camera pose $\mathbf{G}$.
4.2 Comparison with SOTA
Replica dataset. We evaluate on the 8 scenes as Nice-SLAM and iMAP. As shown in Tab. 1, the results demonstrate comparable performance across almost all scenes in terms of Depth L1 and PSNR metrics. Notably, our approach exhibits performance on par with SLAM systems that utilize ground truth depth for supervision.
@@ 4.2 Comparison with SOTA
Table 1: Photometric (PSNR [dB], SSIM, LPIPS) and Geomertric (Depth L1 [cm]) results on Replica dataset [28]. Higher PSNR values mean better color rendering performance and lower Depth L1 values denote better depth reconstruction.
Method Metric Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4 Avg.
Nice-SLAM[52] (GT depth) PSNR $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
SSIM $\uparrow$ 0.63 0.75 0.81 0.87 0.88 0.79 0.85 0.89 0.83
LPIPS $\downarrow$ 0.33 0.27 0.20 0.22 0.18 0.23 0.20 0.19 0.23
Depth L1 $\downarrow$ 3.53 3.60 3.03 5.56 3.35 4.71 3.84 3.35 3.87
Vox-Fusion [44] (GT depth) PSNR $\uparrow$ 22.39 22.36 23.92 27.79 29.83 20.33 23.47 25.21 24.41
SSIM $\uparrow$ 0.68 0.75 0.79 0.85 0.87 0.79 0.80 0.84 0.80
LPIPS $\downarrow$ 0.30 0.26 0.23 0.24 0.18 0.24 0.21 0.19 0.23
Depth L1 $\downarrow$ 2.51 1.69 3.37 2.55 2.08 3.03 2.89 2.86 2.67
SplaTAM [10] (GT depth) PSNR $\uparrow$ 32.86 33.89 35.25 38.26 39.17 31.97 29.70 31.81 34.11
SSIM $\uparrow$ 0.98 0.97 0.98 0.98 0.98 0.97 0.95 0.95 0.97
LPIPS $\downarrow$ 0.07 0.10 0.08 0.09 0.09 0.10 0.12 0.15 0.10
Depth L1 $\downarrow$ 0.63 0.62 0.98 0.57 1.66 7.32 3.94 0.88 2.08
Ours (GT depth) PSNR $\uparrow$ 33.24 34.81 34.16 39.32 39.51 34.08 32.65 34.93 35.34
SSIM $\uparrow$ 0.98 0.99 0.98 0.99 0.99 0.98 0.96 0.98 0.98
LPIPS $\downarrow$ 0.13 0.11 0.11 0.10 0.10 0.13 0.13 0.11 0.12
Depth L1 $\downarrow$ 0.82 0.77 1.35 1.74 1.56 4.23 3.23 1.26 1.87
COLMAP [26] (Predicted depth) PSNR $\uparrow$ 20.93 11.67 10.35 5.88 5.88 15.66 13.73 17.47 12.70
SSIM $\uparrow$ 0.77 0.69 0.75 0.59 0.53 0.80 0.79 0.81 0.72
LPIPS $\downarrow$ 0.29 0.44 0.33 0.44 0.33 0.30 0.29 0.27 0.34
Depth L1 $\downarrow$ 3.87 27.29 5.41 5.21 12.69 4.28 5.29 5.45 8.69
Droid-SLAM [34] (Predicted depth) PSNR $\uparrow$ 18.25 18.65 13.49 16.13 10.31 14.78 15.53 15.71 15.36
SSIM $\uparrow$ 0.73 0.79 0.78 0.76 0.65 0.80 0.79 0.80 0.76
LPIPS $\downarrow$ 0.35 0.28 0.29 0.29 0.28 0.30 0.30 0.31 0.30
Depth L1 $\downarrow$ 12.18 8.35 3.26 3.01 2.39 5.66 4.49 4.65 5.50
Nicer-SLAM[51] (Predicted depth) PSNR $\uparrow$ 25.33 23.92 26.12 28.54 25.86 21.95 26.13 25.47 25.41
SSIM $\uparrow$ 0.75 0.77 0.83 0.86 0.85 0.82 0.85 0.86 0.82
LPIPS $\downarrow$ 0.25 0.21 0.17 0.17 0.17 0.19 0.16 0.17 0.19
Depth L1 $\downarrow$ 2.53 3.93 3.40 5.49 3.45 4.02 3.34 3.03 3.65
Ours (Predicted depth) PSNR $\uparrow$ 29.58 32.74 31.25 36.31 37.22 30.68 30.21 31.96 32.49
SSIM $\uparrow$ 0.83 0.91 0.87 0.94 0.94 0.90 0.88 0.89 0.89
LPIPS $\downarrow$ 0.18 0.16 0.15 0.13 0.15 0.20 0.19 0.18 0.17
Depth L1 $\downarrow$ 1.32 1.19 2.48 3.29 3.07 3.87 4.24 2.78 2.76
360 desk desk2 floor plant room rpy teddy xyz avg
ORB-SLAM2 [18] - 0.071 - 0.023 - - - - 0.010 -
ORB-SLAM3 [1] - 0.017 0.210 - 0.034 - - - 0.009 -
Ours 0.086 0.013 0.023 0.026 0.027 0.049 0.021 0.049 0.009 0.033
Table 3: ATE [m] Results on TUM dataset [29] with RGB-D inputs from freiburg1, freiburg2 and freiburg3 set. Method fr1/desk fr2/xyz fr3/office Kintinuous [19] 0.037 0.029 0.030 BAD-SLAM [27] 0.017 0.011 0.017 ORB-SLAM2 [18] 0.016 0.004 0.010 iMAP [30] 0.049 0.020 0.058 NICE-SLAM [52] 0.027 0.018 0.030 Ours 0.014 0.005 0.011
Table 4: Comparison of runtime and memory usage. Tested on Replica office0 on NVIDIA A6000 GPU. Method Memory Tracking Mapping Usage /frame /frame iMAP [30] 11.72 GB 0.14s 0.32s Nice-SLAM [52] 10.13 GB 1.32s 10.92s Vox-Fusion [44] 12.45 GB 0.36s 0.55s Ours 14.59 GB 0.44s 0.89s
Runtime and Memory Usage. In Table 4, we also report the runtime and memory usage on the Replica dataset. The tracking and mapping time is reported per frame. It can be observed that our method can achieve comparable speed with Vox-Fusion [44], but providing much higher rendering quality as shown in Tab 1.
@@
4.3 Alation Study
4.3.1 Depth Correction.
In Tab. 6 (a), we compare the performance with or without the proposed quadric depth correction module. It can be observed that the depth correction process enhances the accuracy of the predicted depth map by introducing additional surface constraints, rectifying drifting points that deviate from the surfaces. This improvement positively impacts both tracking and mapping performance, especially for the Depth L1 metric.
Table 6: The ablation study of mapping and tracking on Replica dataset office0.
Depth L1 [cm] $\downarrow$ PSNR [dB] $\uparrow$ ATE RMSE [cm] $\downarrow$
(a) Ours (w/o quadric depth correction) 5.63 32.55 0.48
Ours (full, with quadric depth correction) 3.29 36.31 0.39
(b) Ours (w/o quadric ray transformer) 5.02 33.13 0.47
Ours (w only intra-ray transformer) 4.36 34.78 0.44
Ours (full, with quadric ray transformer) 3.29 36.31 0.39
(c) Ours (w/o joint optimization) 4.42 34.21 0.41
Ours (full, with joint optimization) 3.29 36.31 0.39
(d) Ours (w/o semantic supervision) 5.87 31.18 0.51
Ours (full, with semantic supervision) 3.29 36.31 0.39
4.3.2 Quadric Ray Transformer.
In experiments without the quadric ray transformer, the points along a sampled ray are processed independently, lacking feature interaction between each other. The introduction of the intra-ray transformer can help capture information along the ray, and hence improve the performance. As illustrated in Tab. 6 (b), the model incorporating both intra-ray and inter-ray transformer exhibits the best performance, and the inclusion of a single intra-ray transformer also contributes to accuracy improvement.
4.3.3 Joint Optimization.
For experiments without joint optimization, the estimated camera poses are not optimized using the rendering loss during the mapping process. It can be observed from Tab. 6 (c) that joint optimization of camera poses and 3D reconstruction demonstrates a modest improvement in both tracking and mapping accuracy. We believe that the limited improvement stems from the fact that joint optimization is solely conducted on keyframes during the mapping process. Given that keyframes constitute only a small portion of the entire sequence, while the evaluation of camera poses spans all frames, this potentially constrains the impact of joint optimization.
4.3.4 Semantic Supervision.
With an additional semantic head, our model not only renders color and depth but also generates the semantic map. It’s intuitive to utilize segmentation results from previous modules as a supervisory signal, enriching the available information. As shown in Table 6 (d), incorporating semantic supervision can significantly enhance performance.

### v2

@@ 3.4 Joint Optimization of Rendering and Pose Estimation
During mapping, we map the 3D coordinates $\mathbf{x}$, and viewing directions $\mathbf{d}$ to volume density, color and semantic logits $(\sigma,\textit{{c}},\textit{{s}})$. Subsequently, volume rendering is utilized to reconstruct color $\hat{\mathbf{C}}$, depth $\hat{\mathbf{D}}$, and semantic map $\hat{\mathbf{S}}$. For points $\{p_{i}|i=1,\cdots,M\}$ on a fitted quadric surface, the color and depth loss is calculated as follows:
@@
Here, the fitting error $\varepsilon$ serves as an uncertainty penalty. In cases where the fitting is not perfect, the quadric plane is less likely to exhibit similar texture and spatial features, resulting in reduced usefulness of the quadric transformer. Consequently, such patches are down-weighted by incorporating the fitting error $\varepsilon$. Following [38], we use a multi-class cross-entropy loss as the semantic loss $\mathcal{L}_{s}$. Hence the rendering loss is a weighted sum of $\mathcal{L}_{c}$, $\mathcal{L}_{d}$ and $\mathcal{L}_{s}$ with hyperparameters $\lambda_{1},\lambda_{2}$.
@@
For joint optimization, the camera pose $\mathbf{G}=\{g_{t}\}$, a $4\times 4$ transformation matrix, is converted to quaternion ($4\times 1$) and translation ($3\times 1$) vector, which are then taken as trainable parameters to the network, optimized together with the NeRF parameters while minimizing the rendering loss. In this way, the camera poses $\mathbf{G}$ and network parameters $\Theta$ can be jointly optimized.
@@ 4.1 Experimental Setup
Dataset. Q-SLAM is evaluated on a variety of datasets, including Replica [39], ScanNet [40], and KITTI MOT dataset [41]. For evaluation of the reconstruction quality, we test our method on 8 synthetic scenes from Replica, which provides high-quality synthetic scenes. Following GO-SLAM [2], we evaluate the tracking accuracy on ScanNet dataset which offers extensively annotated RGB-D scans of real-world scenarios, encompassing challenging short and long trajectories. KITTI MOT dataset is an outdoor autonomous driving dataset with dynamic objects, such as cars. For camera tracking assessment, our approach is evaluated under two distinct modes: one utilizing ground truth and the other utilizing estimated depth from monocular images as inputs.
@@
Metrics. We evaluate tracking accuracy by aligning the estimated trajectory with the ground truth trajectory and computing the Root Mean Square Error (RMSE) of the Absolute Trajectory Error (ATE). This metric quantifies the Euclidean distance between the estimated pose and the corresponding ground truth pose. Following NeRF-SLAM [4], we utilize Peak Signal-to-Noise Ratio (PSNR), SSIM [42], and LPIPS [43] for image rendering evaluation, which evaluate the similarity between the rendered images and the ground truth images, and Accuracy [cm], Completion [cm], Completion Ratio [%] for 3D reconstruction assessment, which measures the reconstruction quality by comparing the reconstructed mesh and the ground truth mesh.
@@
Table 1: Photometric (PSNR [dB], SSIM, LPIPS) and Geomertric (Acc. [cm], Comp. [cm], Comp. Ratio [%]) results on Replica dataset [39].
Setting Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ATE $\downarrow$ Acc. $\downarrow$ Comp. $\downarrow$ Comp. Ratio (%)  $\uparrow$
RGBD Nice-SLAM [8] 24.42 0.81 0.23 1.95 3.87 3.87 82.41
Vox-Fusion [44] 24.41 0.80 0.24 0.54 2.67 4.55 86.59
SplaTAM [18] 34.11 0.97 0.10 0.36 2.88 3.57 71.68
Ours 35.34 0.98 0.12 0.34 2.17 2.47 91.13
Mono COLMAP [45] 13.94 0.72 0.35 - 8.69 12.12 67.62
DROID-SLAM [20] 21.69 0.81 0.25 0.42 5.50 12.29 63.62
Nicer-SLAM [46] 25.41 0.83 0.19 1.88 3.65 4.16 79.37
Ours 32.49 0.89 0.17 0.38 2.89 3.55 84.79
@@
Implementation Details. All experiments are conducted on an NVIDIA A6000 GPU using PyTorch 1.10.0. We use Adam as the optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The tracking module is adapted from Droid-SLAM [20], and pretrained weights are utilized to estimate depths and poses. Our mapping backbone is adapted from Point-SLAM [17], incorporating our proposed depth correction and quadric ray transformer. For faster image segmentation, we use the MobileSAM [47] to obtain segmentation masks from each input RGB frame. Diverging from methodologies where scene meshes are reconstructed using marching cubes on Signed Distance Function (SDF) values of queried points, Q-SLAM renders images and depths over the estimated camera trajectory and use TSDF-Fusion [36] for mesh construction with voxel size 1 cm. During the fitting process, only quadric surfaces with fitting error lower than the specified threshold undergo depth correction; otherwise, the uncorrected predicted depth is used for the supervision of NeRF training. For joint optimization, the scene representation parameters $\Theta$ are optimized for five steps, and the accumulated losses are then utilized to update camera pose $\mathbf{G}$.
4.2 Comparison with SOTA
Replica dataset. We evaluate on the 8 scenes as Nice-SLAM and iMAP. As shown in Tab. 1, the results demonstrate comparable performance across almost all scenes in terms of rendering and reconstruction metrics. Notably, our approach exhibits performance on par with SLAM systems that utilize ground truth depth for supervision.
Table 3: Runtime and memory usage.
Method Memory Tracking Mapping
Nice-SLAM [8] 10.13 GB 1.32s 10.92s
Vox-Fusion [44] 12.45 GB 0.36s 0.55s
Ours 14.59 GB 0.44s 0.89s
Runtime and Memory Usage. In Table 3, we also report the runtime and memory usage on the Replica dataset. The tracking and mapping time is reported per frame. It can be observed that our method can achieve comparable speed with Vox-Fusion [44], but providing much higher rendering quality as shown in Tab 1. We also report the runtime of individual breakdown components in Tab 5.
4.3 Ablation Study
Depth Correction. In Tab. 5 (a), we compare the performance with or without the proposed quadric depth correction module. It can be observed that the depth correction process enhances the accuracy of the predicted depth map by introducing additional surface constraints, rectifying drifting points that deviate from the surfaces. This improvement positively impacts both tracking and mapping performance, especially for the Depth L1 metric.
Table 4: Ablation on Replica dataset.
Setting Depth L1 PSNR ATE
Full Setting 2.76 32.49 0.38
(a) w/o depth correction 3.23 31.42 0.40
(b) w/o quadric transformer 2.98 31.87 0.39
(c) w/o joint optimization 2.81 32.14 0.42
(d) w/o semantic supervision 2.91 32.09 0.39
Table 5: Breakdown of runtime.
Components Timing (per frame)
Droid-SLAM prediction 54 ms
MobileSAM segmentation 37 ms
Depth Correction 143 ms
Tracking (pose optimization) 442 ms
Mapping 894 ms
@@ 4.3 Ablation Study
Quadric Transformer. In experiments without the quadric ray transformer, the points along a sampled ray are processed independently, lacking feature interaction between each other. The introduction of the quadric ray transformer can help capture information along the ray, and hence improve the performance, as illustrated in Tab. 5 (b).
Joint Optimization. For experiments without joint optimization, the estimated camera poses are not optimized using the rendering loss during the mapping process. It can be observed from Tab. 5 (c) that joint optimization of camera poses and 3D reconstruction demonstrates a modest improvement in both tracking and mapping accuracy. We believe that the limited improvement stems from the fact that joint optimization is solely conducted on keyframes during the mapping process. Given that keyframes constitute only a small portion of the entire sequence, while the evaluation of camera poses spans all frames, this potentially constrains the impact of joint optimization.
Semantic Supervision. With an additional semantic head, our model not only renders color and depth but also generates the semantic map. It’s intuitive to utilize segmentation results from previous modules as a supervisory signal, enriching the available information. As shown in Table 5 (d), incorporating semantic supervision can significantly enhance performance.
@@ 5 Conclusion
Limitations. The heavy computational cost of our method compared to traditional SLAM systems and the difficulty in scaling to unbounded outdoor scenarios. We also acknowledge that the quadric assumption of the surface is not ubiquitous, especially for complex structures in outdoor scenarios, such as trees. Additionally, the smoothing process heavily relies on the segmentation results. We plan to explore more efficient representations, such as gaussian splatting, in the future work.
Acknowledgments
This work is supported by Berkeley DeepDrive.  [^1] https://deepdrive.berkeley.edu
Appendix / Supplemental Materials
Figure 6: Structure overview of Q-SLAM. 1) Tracking: initialize per-frame camera poses and depth prediction. Correct the noisy depth using our proposed depth correction module based on the segmentation results from monocular inputs. 2) NeRF: using the selected keyframes to supervise the optimization of NeRF network equipped with our proposed quadric-decomposed transformer. 3) Mapping: global bundle adjustment to jointly optimize the scene representation and camera poses taking rays sampled from all keyframes. Reconstruct the complete scene by fusing the rendered RGB images and depth maps with TSDF-fusion [36].
Appendix A Additional Methodology Details
A.1 Camera tracking
The initial depth and camera poses are obtained from the learning-based Droid-SLAM [20]. Taking sequences of RGB images as input, Droid-SLAM can predict camera pose and pixel-wise depth through recurrent iterative updates. We do not make changes to the Driod-SLAM codebase but use them to initialize the predicted depths and poses. The rough depths are then rectified with our proposed depth correction module, and the camera poses are optimized with our joint optimization technique.
A.2 Local bundle adjustment
By setting a window size of 25, we will perform a local bundle adjustment. Following Nice-SLAM [8], we treat camera poses as trainable parameters. We sample 400 rays per image within the window and optimize their corresponding poses using the rendering loss.
where $\mathcal{L}_{c}$ is the color loss, $\mathcal{L}_{d}$ is the depth loss and $\mathcal{L}_{s}$ is the semantic loss with hyperparameters $\lambda_{1},\lambda_{2}$.
A.3 Keyframe management
Following Droid-SLAM, our system takes as input a live RGB stream, and applies a recurrent update operator based on RAFT [48] to compute the optical flow of each new frame compared to the last keyframe. If the average flow is larger than a predefined threshold $\tau_{flow}$, a new keyframe is created out of the current frame and added to the maintained keyframe buffer.
A.4 Quadirc surface fitting and depth correction
By setting $\nabla\mathbb{C}_{c}=0$ in Eq. 2 to obtain the optimal $c^{*}$
The cost function in Eq. 2 becomes:
where $\bar{\mathbf{q}}=\frac{1}{N}\sum_{i=1}^{N}q_{i}$ are the quadric term averaged on points in a patch, $\bar{\mathbf{x}}=\frac{1}{N}\sum_{i=1}^{N}x_{i}$ is the linear term.
The intermediate variables are defined as follows:
Setting $\nabla\mathbb{C}_{\mathcal{C}_{l}}=0$ gives
$\mathbb{L}~{}\mathcal{C}_{l}^{*}=\mathbb{N}^{T}\mathcal{C}_{q}$ (12)
By substituting $\mathcal{C}_{l}^{*}$ back to Eq. 10, we can obtain
A.5 Ray points sampling
Initially, rays are constructed from the provided image and calibration matrix, and $N_{s}$ points are uniformly sampled along each ray within the range $[d_{near},d_{far}]$. Utilizing the corrected depth values, denoted as $d$, we refine the sampling process by selecting additional $N_{d}$ samples within the range of $[0.95d,1.05d]$. We do not simply sample around $d$ because potential errors in the corrected depth values might lead to an extended sampling distance away from the true surface.
Appendix B Additional Implementation Details
B.1 Data source
The data of other NeRF-SLAM methods in Tab. 1 is sourced from Nicer-SLAM [46], and the geometric reconstruction results (Acc., Comp., etc.) of SplaTAM [18] comes from RTG-SLAM since the original paper does not report these metrics. The results of other method in Tab. 2 are mainly taken from GO-SLAM [2].
B.2 Dataset
Q-SLAM is evaluated on a variety of datasets, including Replica [39], ScanNet [40], and TUM RGB-D [50] dataset. For evaluation of reconstruction quality, we test our method on 8 synthetic scenes from Replica, which provides high-quality synthetic scenes, akin to the evaluation framework adopted by NeRF-SLAM [4]. Following GO-SLAM [2], we evaluate the tracking accuracy on ScanNet dataset which offers extensively annotated RGB-D scans of real-world scenarios, encompassing challenging short and long trajectories. Following Nice-SLAM [8], we also evaluate on various scenes on indoor TUM RGB-D dataset, with ground truth poses provided by a motion capture system. For camera tracking assessment, our approach is tested under two distinct modes: one utilizing ground truth and the other utilizing estimated depth from monocular images as inputs. The batch size of sampled rays to NeRF network is 8192.
B.3 Evaluation Metrics
We evaluate tracking accuracy by aligning the estimated trajectory with the ground truth trajectory and computing the Root Mean Square Error (RMSE) of the Absolute Trajectory Error (ATE). This metric quantifies the Euclidean distance between the estimated pose and the corresponding ground truth pose. In line with the evaluation approach of NeRF-SLAM [4], we utilize Peak Signal-to-Noise Ratio (PSNR), SSIM [42], and LPIPS [43] for image rendering evaluation, and Accuracy [cm], Completion [cm], Completion Ratio [%] for 3D reconstruction assessment.
  • Absolute Trajectory Error (ATE) (cm) $\downarrow$: Evaluates trajectory estimation accuracy by measuring the average Euclidean translation distance between corresponding poses in estimated and ground truth trajectories, often reported in terms of Root Mean Square Error (RMSE).
  • Peak Signal to Noise Ratio (PSNR) $\uparrow$: Measures image quality by evaluating the ratio between the maximum pixel value and the root mean squared error, usually expressed in terms of the logarithmic decibel scale.
  • Structural Similarity Index Measure (SSIM) $\uparrow$: Assesses image quality by examining the similarities in luminance, contrast, and structural information among patches of pixels.
  • Learned Perceptual Image Patch Similarity (LPIPS) $\downarrow$: Utilizes learned convolutional features to assess image quality based on feature map mean squared error across layers.
  • Accuracy (cm) $\downarrow$: Computes the average distance between sampled points from the reconstructed mesh and the nearest ground-truth point.
  • Completion (cm) $\downarrow$: Measures the average distance between sampled points from the ground-truth mesh and the nearest reconstructed.
  • Completion Ratio (%) $\uparrow$: the percentage of points in the reconstructed mesh with Completion under 5 cm.
B.4 Hyperparameters
All experiments are conducted on NVIDIA A6000 GPU with PyTorch 1.10.0. We use Adam as our optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$. The tracking backbone is Droid-SLAM, where we use the pretrained weights to estimate depths and poses. We use Point-SLAM as the mapping backbone, equipped with our proposed depth correction and quadric transformer. The threshold for motion filter is 4.0 pixels, a tracked frame is considered as a keyframe only if the average optical flow is greater than the threshold. The window size for local bundle adjustment is 25. During the joint optimization process, camera poses are optimized for one epoch, and the NeRF network parameters are optimized for five epochs.
B.5 Segmentation and Quadric Fitting
For the segmentation network, we use Segment Anything Model (SAM) [51] and MobileSAM [47] for faster inference, an off-the-shelf network to produce the mask for quadric fitting. To prevent the negative effect of outliers, quadric fitting only applies to segments with area larger than 200 pixels.
During the fitting process, we calculate the coefficient of determination to evaluate the fitting performance. Let $z_{i}$ be the predicted depth value and $f_{i}$ be the corresponding corrected depth. The coefficient of determination is calculated as follows:
$\bar{z}$ $=\frac{1}{n}\sum_{i=1}^{n}z_{i}$
$SS_{\mathrm{res}}$ $=\sum_{i}\left(z_{i}-f_{i}\right)^{2}$
$SS_{\mathrm{tot}}$ $=\sum_{i}\left(z_{i}-\bar{z}\right)^{2}$
$R^{2}$ $=1-\frac{SS_{\mathrm{res}}}{SS_{\mathrm{tot}}}$
We only perform depth correction on quadric surfaces with fitting coefficient greater than the given threshold 0.85, otherwise, we just use the predicted depth for the supervision of NeRF training.
B.6 Mesh Reconstruction
Different from other approaches that reconstruct mesh of a scene by running marching cubes on the Signed Distance Function (SDF) values of the queried points, we render images and depths for the selected keyframes. The reason for the difference is that our rendering requires to correlate points along and across rays, while other approaches process 3D points independently. We first render RGB images and depth maps, and then use TSDF-fusion [36] to reconstruct the 3D volume mesh.
Appendix C Additional Experimental Results
360 desk desk2 floor plant room rpy teddy xyz avg
ORB-SLAM2 [52] - 0.071 - 0.023 - - - - 0.010 -
ORB-SLAM3 [53] - 0.017 0.210 - 0.034 - - - 0.009 -
Ours 0.086 0.013 0.023 0.026 0.027 0.049 0.021 0.049 0.009 0.033
Table 7: ATE [m] Results on TUM dataset [50] with RGB-D inputs from freiburg1, freiburg2 and freiburg3 set.
Method fr1/desk fr2/xyz fr3/office
Kintinuous [56] 0.037 0.029 0.030
BAD-SLAM [57] 0.017 0.011 0.017
ORB-SLAM2 [52] 0.016 0.004 0.010
iMAP [5] 0.049 0.020 0.058
NICE-SLAM [8] 0.027 0.018 0.030
Ours 0.014 0.005 0.011
Appendix D Visualization
In Fig. 7, we provide the qualitative results of the reconstruction. It can be observed that our method outperforms GO-SLAM, especially on the boundaries of objects.
We provide several selected visualization results for depth correction as shown in Fig. 8. Benefiting from the segmentation mask, the depth correction improves the sharpness of the boundary of objects.
Figure 8: Qualitative results of depth correction.